Find the inverse Laplace transform of the form of the given function. A.) 3S g²-5-6 B.) Les-3 S²_4
Answer #1

0 58 2 9-SS (or(-3) - 77 { is) T SS 6 2 t 9 q ezt, 2 .e 5 5 5 ALL + E 4S-3 s1 4 (s +) 413-9 117 1 5 + 5-2 S+2 2 TA 4S-3 54 5

\large \frac{6s-3}{s^2-4}:\quad \frac{15}{4\left(s+2\right)}+\frac{9}{4\left(s-2\right)}\ \\ \Rightarrow L^{-1}\left\{\frac{6s-3}{s^2-4}\right\}=L^{-1}\left\{\frac{15}{4\left(s+2\right)}\right\}+L^{-1}\left\{\frac{9}{4\left(s-2\right)}\right\} \\= \frac{15}{4}e^{-2t}+\frac{9}{4}e^{2t} \\ Therefore \\\mathrm{Laplace\:Inverse\:Transform\:of\:}\frac{6s-3}{s^2-4}:\quad \frac{15}{4}e^{-2t}+\frac{9}{4}e^{2t}